# FlowPolicy: Enabling Fast and Robust 3D Flow-based Policy via Consistency Flow Matching for Robot Manipulation

### v1

@@ Abstract
Robots can acquire complex manipulation skills by learning policies from expert demonstrations, which is often known as vision-based imitation learning. Generating policies based on diffusion and flow matching models has been shown to be effective, particularly in robotics manipulation tasks. However, recursion-based approaches are often inference inefficient in working from noise distributions to policy distributions, posing a challenging trade-off between efficiency and quality. This motivates us to propose FlowPolicy, a novel framework for fast policy generation based on consistency flow matching and 3D vision. Our approach refines the flow dynamics by normalizing the self-consistency of the velocity field, enabling the model to derive task execution policies in a single inference step. Specifically, FlowPolicy conditions on the observed 3D point cloud, where consistency flow matching directly defines straight-line flows from different time states to the same action space, while simultaneously constraining their velocity values, that is, we approximate the trajectories from noise to robot actions by normalizing the self-consistency of the velocity field within the action space, thus improving the inference efficiency. We validate the effectiveness of FlowPolicy on Adroit and Metaworld, demonstrating a 7$\times$ increase in inference speed while maintaining competitive average success rates compared to state-of-the-art policy models. Codes will be made publicly available.
@@ Introduction
Enabling robots to acquire diverse manipulation tasks by learning from expert demonstrations, commonly referred to as visual imitation learning, has been a long-standing objective in robotic learning and embodied AI (Atkeson and Schaal 1997; Argall et al. 2009; Wang et al. 2024a; Zare et al. 2024; Li et al. 2024). Prior work attempts to tackle this problem either by explicitly defining it as a supervised regression task (Zhang et al. 2018; Florence, Manuelli, and Tedrake 2019; Toyer et al. 2020; Rahmatizadeh et al. 2018), or by implicitly modeling action distributions using energy-based models (Florence et al. 2022; Jarrett, Bica, and van der Schaar 2020). The former learns the mapping from observations to actions, while the latter seeks to identify the minimal energy action through energy optimization. However, explicit regression approaches are often inadequate for capturing multi-modal demonstration behaviors and struggle with high-precision tasks (Florence et al. 2022). Conversely, energy-based models face challenges with training stability, primarily due to the necessity of negative sample extraction during the training process (Chi et al. 2023).
Figure 1: Comparison of FlowPolicy with the state-of-the-art method DP3 (Ze et al. 2024) and its lightweight version Simple DP3 in terms of inference time and average success rate on Adroit and Metaworld. Our FlowPolicy shows the best trade-off, achieving competitive results with an average inference time of 19.9ms. FlowPolicy is 7$\times$ faster than DP3 (19.9ms vs. 145.7ms) and 3$\times$ faster than Simple DP3 (19.9ms vs. 63.0ms).
@@
On the other hand, Flow Matching (FM) (Lipman et al. 2022), another class of generative model frameworks that directly defines probability paths via Ordinary Differential Equations (ODEs) to transform between noise and data samples, has demonstrated greater numerical stability and inference efficiency than diffusion families. The latest advancement in this area, Consistency Flow Matching (CFM) (Yang et al. 2024), further improves efficiency by directly defining a straight-line flow from any point in time to the same endpoint, enabling one-step generation for FM, which offers a potential direction for real-time policy generation. However, the conditional generation capability of CFM, particularly for embodied tasks involving complex 3D vision representations as conditions, remains underexplored.
In this paper, we address these challenges in policy generation by leveraging the concept of consistency flow matching, introducing a novel 3D flow-based framework for real-time robotic manipulation policy generation, named FlowPolicy. More specifically, we make the first attempt at conditional consistent flow matching in 3D robotic manipulation tasks. FlowPolicy allows for the fastest possible generation of target actions by directly defining straight-line flow conditions on the initial 3D point clouds. In addition, multi-segment training is applied to ensure the quality of action generation, aiming at a better trade-off between efficiency and effectiveness. We evaluate our framework on 37 robot manipulation tasks from Adroit and Metaworld, encompassing a wide range of tasks from simple to complex and dexterous manipulation. As illustrated in Fig. 1, our approach achieves a 7$\times$ reduction in average inference time while maintaining a competitive average success rate compared to state-of-the-art methods based on 3D visual conditional representation. By combining consistency flow matching with 3D condition-based robot manipulation, We enable the real-time decoding of high-quality robot actions with on-step inference.
Figure 2: Overall pipeline. The above section is a visualization of FlowPolicy. By defining a straight-line flow the data can flow the fastest from the noise distribution to the action distribution (Adroit: Open the door). The following section shows the details of FlowPolicy. Given a certain number of expert presentations, it is first converted into 3D point clouds. The 3D point clouds and the robot state are then fed into two encoders to obtain the compact 3D visual representation and the robot state embedding, respectively. Finally, a straight-line flow is learned by conditional consistency flow matching to generate high-quality robot actions and perform the corresponding tasks (Metaworld: Assembly) at real-time inference speed.
@@
Diffusion models are a family of generative models that are highly generalizable and easy to train, which are initially utilized for image generation tasks (Batzolis et al. 2021; Epstein et al. 2023). As a subversive approach, diffusion models progressively refine Gaussian noise by recursively refining it to the point where it satisfies the data distribution. Recently, it has been introduced to the field of robotics, including reinforcement learning (Römer et al. 2024), imitation learning (Wang et al. 2024b), trajectory Optimization (Carvalho et al. 2023) and grasping (Urain et al. 2023), and has shown impressive performance. Denoising Diffusion Probabilistic Models (DDPM) (Ho, Jain, and Abbeel 2020) generates a clean sample from noise by solving a Stochastic Differential Equation (SDE). However, numerical simulations of diffusion processes typically involve hundreds of steps, leading to significant slowness. Denoising Diffusion Implicit Modeling (DDIM) (Song, Meng, and Ermon 2020) speeds up inference by solving the inverse process as an Ordinary Differential Equation (ODE), but this sacrifices generation quality. Although adaptive ODE solvers exist, they do not significantly reduce the number of inference steps. In 3D-related robot manipulation tasks with high-dimensional control, it is challenging to achieve real-time inference speed with high-quality actions based on diffusion models. The main obstacles arise from the uncertainty of the diffusion process and its recursive nature.
To improve the efficiency problem of the diffusion process, Song (Song et al. 2023) proposes the consistency model. The consistency model is a novel model that directly maps noise to a data distribution by learning to map points at arbitrary moments on the probabilistic flow ODE (PF-ODE) trajectory to the origin (i.e., clean data) to support one-step or several-step generation. It has been utilized in several domains such as high fidelity image generation (Luo et al. 2023), text (Seo et al. 2023) and video (Chai et al. 2023) due to its efficient expressiveness. While consistency models support one-step generation, distillation is usually required to reach the desired results. This results in its limited application in robotic manipulations.
@@
Recently, Yang (Yang et al. 2024) proposed to perform consistency in velocity field space rather than sample space to improve the flow matching process (CFM). CFM directly defines a straight-line flow from any time to the same endpoint and imposes constraints on its velocity values. Also, the multi-segment training technique allows CFM to achieve a better trade-off between sampling quality and velocity. Therefore, CFM is an elegant balance of efficiency and performance. More importantly, Yang  (Yang et al. 2024) demonstrated that CFM can be trained to produce a robust flow model without the aid of distillation, which is valuable to robots performing unseen tasks, as it is difficult to find a robust teacher model in practical applications. While consistency-based models have shown promising results in some areas of embodied intelligence (Ding and Jin 2024; Chen, Li, and Zhao 2024; Prasad et al. 2024), its effectiveness has not been demonstrated on embodied tasks with complex 3D vision representation as conditions, especially for flow-based methods. To address this issue, we propose FlowPolicy, a real-time 3D policy generation framework based on consistency flow matching.
@@
We focus on imitation learning to enable robots to acquire complex manipulation skills. Specifically, we expect a limited number of expert demonstrations to teach an agent to learn a policy $\pi:O\Longrightarrow A$, i.e., mapping actions from visual observations $\textit{o}\in{O}$ to actions $\textit{a}\in{A}$. Visual observations include the robot state and scene point clouds, and actions are usually sequences of trajectories of the robot to accomplish a specific task. Diffusion-based and flow-based methods can generate high-quality robot actions, but expensive iterative sampling makes them unable to accomplish real-time operations, and thus the application scenarios are very limited. While CFM balances efficiency and effectiveness, few existing works have explored the application of 3D condition generation based on consistent flow matching to robotic manipulation. Therefore, we propose FlowPolicy, a conditional consistency flow matching model, which guarantees the generation of high-quality actions while also accomplishing one-step inference for real-time applications. This section is organized as follows, we first review flow matching and consistent flow matching, and then introduce the pipeline of FlowPolicy. Finally, we describe the design details of each component.
@@
Flow matching is a simple and attractive objective. However, lack of a prior knowledge about $u$ and $p_{t}$, conditional flow matching(Lipman et al. 2022) is typically employed to regress ${\nu}_{\theta}(t,x_{t})$ given a conditional vector field $u(t,x_{t}|{x_{1}})$ and a conditional probability path $p(t,x_{t}|{x_{1}})$:
$L_{\text{FM}^{\ast}}(\theta)=E_{t,q_{(}{x_{1}})}E_{t,p_{t}|{x_{1}}}{\|{{\nu}_{\theta}(t,x_{t})-u(t,x_{t}|x_{1})}\|}^{2}_{2}$ (3)
Learning straight-line flows enables faster inference efficiency. Consistency flow matching (CFM)  (Yang et al. 2024) is a generalized method for efficiently learning straight-line flows without approximating the entire probabilistic path. Consistent velocity is defined as ${\nu}(t,{\xi}_{x}(t))={\nu}(0,x)$, indicating that the velocity along the solution of Equation 1 remains constant. Due to the complexity of the target distribution solution, CFM does not regress directly on the ground truth vector field, instead, it directly defines a velocity-consistent straight-line flow that departs at different times to the same endpoint. CFM can regress on the following objective to learn a velocity-consistent vector field:
${L_{\theta}=E_{t{\sim}\mu}E_{x_{t},x_{t}+\Delta{t}}\|f_{\theta}(t,x_{t})-f_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}_{2}}\\ {+\alpha{\|{\nu}_{\theta}(t,x_{t})-{\nu}_{\theta^{-}}(t+\Delta{t})\|^{2}_{2}}}\\ f_{\theta}(t,x_{t})=x_{t}+(1-t)*{\nu}_{\theta}(t,x_{t})$ (4)
where $\mu$ is a uniform distribution on $[0,1-\Delta{t}]$, $\alpha$ is a positive scalar, and $\Delta{t}$ is the time interval, a small and positive scalar. ${\theta}^{-}$ denotes the running average of past $\theta$ values computed using an exponential moving average (EMA). Training $L_{\theta}$ not only regularizes the velocity but also learns the data distribution.
Additionally, CFM allows for multi-segment training to enhance its ability to express the distribution of transmission through $p_{t}$. Specifically, multi-segment CFM splits time into uniform subsets, normalizes a consistent vector field ${\nu}(t,x_{t})$ within each subset, and in turn constructs segmented linear trajectories. This approach relaxes the velocity consistency requirement and makes data transmission more flexible and efficient. Given the number of segments K, the training loss of segmented CFM is as follows:
$L_{\theta}=E_{t{\sim}{\mu}^{i}}{\lambda}^{i}E_{x_{t},x_{t}+\Delta{t}}\|f^{i}_{\theta}(t,x_{t})\\ -f^{i}_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}{{}_{2}}\\ {+\alpha{\|{\nu}^{i}_{\theta}(t,x_{t})-{\nu}^{i}_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}{{}_{2}}}}\\ f^{i}_{\theta}(t,x_{t})=x_{t}+(i+1)/k-t*{\nu}^{i}_{\theta}(t,x_{t})$ (5)
@@
Table 1: Quantitative comparison of runtime between state-of-the-art policy models and the proposed FlowPolicy. We evaluate 3 dexterity tasks from Adroit and 34 challenging tasks from Metaworld across 3 random seeds and report inference time per step (ms) with standard deviation. Our FlowPolicy enables real-time robot operations with an average time of 19.9ms in one-step inference, which is 7 $\times$ faster than DP3 and 3 $\times$ faster than its lightweight version, SimpleDP3.
@@
Table 2: Comparisons on success rate between state-of-the-art policy models and the proposed FlowPolicy. We evaluate 3 dexterity tasks from Adroit and 34 challenging tasks from Metaworld across 3 random seeds and report the success rate (%) with standard deviation. For fair comparisons, DP3 and Simple DP3 are reproduced under the same experimental environment and expert demonstrations as ours. Our method achieves competitive success rates compared to the current state-of-the-art approach DP3 while enabling real-time, one-step inference.
@@
FlowPolicy applies 3D point clouds as a guide to generate execution-specific policies given its rich spatial information. Specifically, we first acquire single-view depth-observation images and then convert the depth information into a point cloud via camera intrinsics and extrinsic. The single-view image is more practically applicable and the point cloud provides a richer spatial representation. Further, a farthest point sampling (FPS) (Qi et al. 2017) is applied to downsample the point cloud to preserve spatial wholeness and reduce redundancy. Finally, we employ a lightweight MLP to encode the point cloud into a compact 3D representation as the visual conditional representation for flow-based policy. With the conditional constraints of the point cloud data, FlowPolicy enables the efficient capture of spatial information that is critical to fine robotic manipulation tasks.
@@
To train a high-quality one-step generator, we employ a two-segment CFM and use backpropagation to minimize the following equation:
@@
By training the two-segment loss above two consistent vector fields can be learned, which in turn constructs a two-segment linear trajectory that transforms the noisy trajectory into the target trajectory. In the inference process, the noise trajectories are sampled from the source distribution $a_{0}\in a^{\text{src}}$ and then predicted by a flow model to obtain action trajectories $a_{1}\in a^{\text{tar}}$.
FlowPolicy provides a framework for generating high-quality actions in a one-step process that promotes a trade-off between efficiency and effectiveness. Compared to DP3 and Simple DP3, FlowPolicy utilizes minimal inference time to achieve a competitive average success rate.
@@
We have chosen two preeminent environmental simulators, Adroit (Rajeswaran et al. 2018) and Metaworld (Yu et al. 2021), as our benchmarks for a comprehensive set of 37 tasks. The Adroit benchmark presents three challenges of dexterous manipulation tasks, where the execution of complex tasks is facilitated by the intricate control of a high-dimensional, articulated hand. In contrast, the Metaworld benchmark offers a diverse array of tasks that span the spectrum of difficulty levels, from easy to very hard, typically surmounted with the aid of a two-fingered robotic arm. For the procurement of expert demonstrations, we have harnessed a well-trained heuristic strategy, designed to yield datasets of exceptional quality.
@@
The main objective of this work is to leverage 3D point clouds as visual conditional representations, aiming to elevate the robot manipulation to the real-time level via a consistency flow matching strategy. Therefore, we focus on comparing DP3, a state-of-the-art 3D conditional diffusion-based strategy model. In addition, its lightweight version, simple DP3, is also compared. For a fair comparison, the DP3 and Simple DP3 parameter settings in our experimental setup are the same as in (Ze et al. 2024).
@@
We evaluate 20 episodes per 200 training epochs and record the mean of the five highest success rates. To avoid chance error, each task is run repeatedly under three different random seeds and their means and variances as well as inference times are calculated. In addition, the Number of Function Evaluations (NFE) (Yang et al. 2024) is evaluated as a metric of inference efficiency. The higher the success rate, the better, and the lower the NFE and inference time, the better.
@@
For each task, we generate a compact conditional embedding consisting of a 3D point cloud and the robot’s state with a dimension of 64. Additionally, our model employs an observation horizon of two steps, signifying that it leverages the point clouds from the two most recent time frames as conditional inputs for policy generation. Focused on efficient imitation learning, we assign ten expert demonstrations to each task. To mitigate unnecessary computational costs, the image size is cropped to a resolution of 84$\times$84 pixels and the point cloud is downsampled to 512 points for Adroit or 1024 points for Metaworld using the farthest point sampling method. During the training phase, the weights are updated using the AdamW optimizer with a learning rate of 1e-4, and a batch size of 128. The training process is iterated for 3,000 epochs. EMA decay rate is set to the default value of 0.95. Additionally, both action and state inputs are normalized to the interval [-1, 1] to facilitate the training process, with actions being unnormalized before execution. Our model and all compared methods are developed based on the PyTorch framework and evaluated on a single NVIDIA RTX 2080Ti GPU.
@@
Table 1 reports the average inference time of each model for 3 dexterity tasks from Adroit and 34 different difficulty tasks from Metaworld. Our FlowPolicy shows impressive results by reducing the average inference time to 19.9ms, an improvement of more than 7$\times$ (19.9ms vs. 145.7ms) compared to the DP3 model. Even for the lightweight version of Simple DP3, FlowPolicy improves the inference efficiency by more than 3$\times$ (19.9ms vs. 63.0ms). Since a consistent velocity field provides a more direct linear path result in generating actions faster and more efficiently. More interestingly, our approach has less variation in inference time. Specifically, FlowPolicy has less fluctuation in the average runtime of performing individual tasks and overall tasks, with a potential explanation attributed to the deterministic sampling policy of consistent flow matching. The results show that faster ODE trajectories can be determined and the inference efficiency can be improved by directly defining a straight-line flow, which is crucial for realizing real-time robot operation.

### v2

@@ Abstract
Robots can acquire complex manipulation skills by learning policies from expert demonstrations, which is often known as vision-based imitation learning. Generating policies based on diffusion and flow matching models has been shown to be effective, particularly in robotic manipulation tasks. However, recursion-based approaches are inference inefficient in working from noise distributions to policy distributions, posing a challenging trade-off between efficiency and quality. This motivates us to propose FlowPolicy, a novel framework for fast policy generation based on consistency flow matching and 3D vision. Our approach refines the flow dynamics by normalizing the self-consistency of the velocity field, enabling the model to derive task execution policies in a single inference step. Specifically, FlowPolicy conditions on the observed 3D point cloud, where consistency flow matching directly defines straight-line flows from different time states to the same action space, while simultaneously constraining their velocity values, that is, we approximate the trajectories from noise to robot actions by normalizing the self-consistency of the velocity field within the action space, thus improving the inference efficiency. We validate the effectiveness of FlowPolicy in Adroit and Metaworld, demonstrating a 7$\times$ increase in inference speed while maintaining competitive average success rates compared to state-of-the-art methods. Code is available at https://github.com/zql-kk/FlowPolicy.
@@ Introduction
Enabling robots to acquire diverse manipulation tasks by learning from expert demonstrations, commonly referred to as imitation learning, has been a long-standing objective in robotic learning and embodied AI (Atkeson and Schaal 1997; Argall et al. 2009; Wang et al. 2024a; Zare et al. 2024; Li et al. 2024). Prior work attempts to tackle this problem either by explicitly defining it as a supervised regression task (Zhang et al. 2018; Florence, Manuelli, and Tedrake 2019; Toyer et al. 2020; Rahmatizadeh et al. 2018), or by implicitly modeling action distributions using energy-based models (Florence et al. 2022; Jarrett, Bica, and van der Schaar 2020). The former learns the mapping from observations to actions, while the latter seeks to identify the minimal energy action through energy optimization. However, explicit regression approaches are often inadequate for capturing multi-modal demonstration behaviors and struggle with high-precision tasks (Florence et al. 2022). Conversely, energy-based models face challenges with training stability, primarily due to the necessity of negative sample extraction during the training process (Chi et al. 2023).
Figure 1: Comparison of FlowPolicy with the state-of-the-art 2D-based method DP(Chi et al. 2023) and 3D-based methods DP3 (Ze et al. 2024) and its lightweight version Simple DP3 in terms of inference time and average success rate on Adroit and Metaworld.
@@
On the other hand, Flow Matching (FM) (Lipman et al. 2022), another class of generative model frameworks that directly defines probability paths via Ordinary Differential Equations (ODEs) to transform between noise and data samples, has shown greater numerical stability and inference efficiency than diffusion families. The latest advancement in this area, Consistency Flow Matching (Consistency-FM) (Yang et al. 2024), further improves efficiency by defining a straight-line flow from any point in time to the same endpoint, making one-step generation possible, which offers a potential direction for real-time policy generation. However, the conditional generation capability of Consistency-FM, particularly for embodied tasks involving complex 3D vision representations as conditions, remains underexplored.
In this paper, we address these challenges in policy generation by leveraging the concept of consistency flow matching, introducing a novel 3D flow-based framework for real-time robotic manipulation policy generation, named FlowPolicy. More specifically, we make the first attempt at conditional consistent flow matching in 3D robotic manipulation tasks. FlowPolicy allows for the fastest possible generation of target actions by directly defining straight-line flow conditions on the initial 3D point clouds. In addition, multi-segment training is applied to ensure the quality of action generation, aiming at a better trade-off between efficiency and effectiveness. We evaluated our framework on 37 robot manipulation tasks from Adroit and Metaworld. As illustrated in Fig. 1, our approach achieves a 7$\times$ reduction in average inference time while maintaining a competitive average success rate compared to state-of-the-art methods based on 3D visual conditional representation. By combining consistency flow matching with 3D condition-based robot manipulation, we enable the real-time decoding of high-quality robot actions with one-step inference.
Figure 2: Overall pipeline. The top section visualizes FlowPolicy, where a straight-line flow enables the fastest data transition from the noise distribution to the action distribution (Adroit: Open the door). The bottom section shows the details of FlowPolicy: expert demonstrations are converted to 3D point clouds, which, along with the robot state, are encoded into compact 3D visual representations and state embeddings. A straight-line flow is then learned via conditional consistency flow matching, generating high-quality actions for tasks (Metaworld: Assembly) at real-time inference speed.
@@
Diffusion models are a family of generative models that are highly generalizable and easy to train, which are initially utilized for image generation tasks (Batzolis et al. 2021; Epstein et al. 2023). As a subversive approach, diffusion models progressively refine Gaussian noise by recursively refining it to the point where it satisfies the data distribution. Recently, it has been introduced to the field of robotics, including reinforcement learning (Römer et al. 2024), imitation learning (Wang et al. 2024b), trajectory optimization (Carvalho et al. 2023) and grasping (Urain et al. 2023), and have shown impressive performance. Denoising Diffusion Probabilistic Models (DDPMs) (Ho, Jain, and Abbeel 2020) generate a clean sample from noise by solving a Stochastic Differential Equation (SDE). However, numerical simulations of diffusion processes typically involve hundreds of steps, greatly slowing the inference speed. Denoising Diffusion Implicit Models (DDIMs) (Song, Meng, and Ermon 2020) speeds up inference by solving the inverse process as an Ordinary Differential Equation (ODE), but this sacrifices generation quality. Although adaptive ODE solvers are available, they do not meaningfully reduce the number of inference steps. In 3D robot manipulation tasks involving high-dimensional control, achieving real-time inference speed with high-quality actions using diffusion models is challenging. The primary obstacles arise from the inherent uncertainty of the diffusion process and its recursive nature.
To improve the efficiency of the diffusion process, Song (Song et al. 2023) proposes the consistency model. The consistency model directly maps noise to a data distribution by learning to map points at arbitrary moments on the probabilistic flow ODE (PF-ODE) trajectory to the origin (i.e., clean data) to support one-step or several-step generation. It has been utilized in several domains such as high-fidelity image generation (Luo et al. 2023), text-to-3D generation (Seo et al. 2023), and video editing (Chai et al. 2023). Although consistency models support one-step generation (Lu et al. 2024), they typically require distillation to achieve the desired results, which restricts their applicability in real-world robotics. Compared to the above methods, our FlowPolicy does not require numerical simulation or distillation.
@@
Recently, Yang (Yang et al. 2024) proposed to perform consistency in velocity field space rather than sample space to improve the flow matching process (Consistency-FM). Consistency-FM directly defines a straight-line flow from any time to the same endpoint and imposes constraints on its velocity values. Also, the multi-segment training technique allows Consistency-FM to achieve a better trade-off between sampling quality and efficiency. Therefore, Consistency-FM is an elegant balance of efficiency and performance. More importantly, Consistency-FM can be trained to produce a robust flow model without the aid of distillation, which is valuable to robots performing unseen tasks, as it is difficult to find a robust teacher model in practical applications. While consistency-based models have shown promising results in some areas of embodied intelligence (Ding and Jin 2024; Chen, Li, and Zhao 2024; Prasad et al. 2024), its effectiveness has not been demonstrated on embodied tasks with complex 3D vision representation as conditions, especially for flow-based methods. To address this issue, we propose FlowPolicy, a real-time 3D policy generation framework based on consistency flow matching.
@@
Our method expects a limited number of expert demonstrations to teach an agent to learn a policy $\pi:O\Longrightarrow A$, i.e., mapping from visual observations $\textit{o}\in{O}$ to actions $\textit{a}\in{A}$. Visual observations include the robot state and scene point clouds, and actions are usually sequences of trajectories of the robot to accomplish a specific task. Diffusion-based and flow-based methods can generate high-quality robot actions, but expensive iterative sampling makes them unable to accomplish real-time operations, and thus the application scenarios are very limited. While Consistency-FM balances efficiency and effectiveness, existing works have not explored the application of 3D condition generation based on consistent flow matching to robotic manipulation. Therefore, we propose FlowPolicy, a conditional consistency flow matching model, which guarantees the generation of high-quality actions while also accomplishing one-step inference for real-time applications. This section is organized as follows: we first review flow matching and consistent flow matching, and then introduce the pipeline of FlowPolicy. Finally, we describe the design details of each component.
@@
Flow matching is a simple and attractive objective. However, lack of a prior knowledge about $u$ and $p_{t}$, conditional flow matching (Lipman et al. 2022) is typically employed to regress ${\nu}_{\theta}(t,x_{t})$ given a conditional vector field $u(t,x_{t}|{x_{1}})$ and a conditional probability path $p_{t}(x_{t}|{x_{1}})$:
$L_{\text{FM}^{\ast}}(\theta)=E_{t,q_{({x_{1}})}}E_{t,{(p_{t}|{x_{1}}})}{\|{{\nu}_{\theta}(t,x_{t})-u(t,x_{t}|x_{1})}\|}^{2}_{2}$ (3)
Where $x_{1}{\sim}q(x_{1})$. Learning straight-line flows enables faster inference efficiency. Consistency flow matching (Consistency-FM)  (Yang et al. 2024) is a generalized method for efficiently learning straight-line flows without approximating the entire probabilistic path. Consistent velocity is defined as ${\nu}(t,{\xi}_{x}(t))={\nu}(0,x)$, indicating that the velocity along the solution of Equation 1 remains constant. Due to the complexity of the target distribution solution, Consistency-FM does not regress directly on the ground truth vector field, instead, it directly defines a velocity-consistent straight-line flow that departs at different times to the same endpoint. Consistency-FM can regress on the following objective to learn a velocity-consistent vector field:
${L({\theta})=E_{t{\sim}\mu}E_{x_{t},x_{t}+\Delta{t}}\|f_{\theta}(t,x_{t})-f_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}_{2}}\\ {+\alpha{\|{\nu}_{\theta}(t,x_{t})-{\nu}_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}_{2}}}\\ f_{\theta}(t,x_{t})=x_{t}+(1-t)*{\nu}_{\theta}(t,x_{t})$ (4)
where $\mu$ is a uniform distribution on $[0,1-\Delta{t}]$, $\alpha$ is a positive scalar, and $\Delta{t}$ is the time interval, a small and positive scalar. ${\theta}^{-}$ denotes the running average of past $\theta$ values computed using an exponential moving average (EMA). Training $L(\theta)$ not only regularizes the velocity but also learns the data distribution.
Additionally, Consistency-FM allows for multi-segment training to enhance its ability to express the distribution of transmission through $p_{t}$. Specifically, multi-segment Consistency-FM splits time into uniform subsets, normalizes a consistent vector field ${\nu}^{i}_{\theta}{(t,x_{t})}$ within each subset, and in turn constructs segmented linear trajectories. This approach relaxes the velocity consistency requirement and makes data transmission more flexible and efficient. Given the number of segments $K$, the training loss of segmented Consistency-FM is as follows:
$L({{\theta}^{\ast}})=E_{t{\sim}{\mu}^{i}}{\lambda}^{i}E_{x_{t},x_{t}+\Delta{t}}\|f^{i}_{\theta}(t,x_{t})\\ -f^{i}_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}_{2}\\ {+\alpha{\|{\nu}^{i}_{\theta}(t,x_{t})-{\nu}^{i}_{\theta^{-}}(t+\Delta{t},x_{t+\Delta{t}})\|^{2}_{2}}}\\ f^{i}_{\theta}(t,x_{t})=x_{t}+((i+1)/k-t)*{\nu}^{i}_{\theta}(t,x_{t})$ (5)
@@
Table 1: Quantitative comparison of runtime between state-of-the-art policy models (Chi et al. 2023; Ze et al. 2024) and the proposed FlowPolicy. We evaluate 37 tasks from Adroit and Metaworld across 3 random seeds and report inference time per step (ms) with standard deviation. Our FlowPolicy enables real-time robot operations with an average time of 19.9ms in one-step inference, which is 7 $\times$ faster than DP3 and 3 $\times$ faster than its lightweight version, SimpleDP3.
@@
Table 2: Comparisons on success rate between state-of-the-art 2D-based (Chi et al. 2023; Hu et al. 2024; Prasad et al. 2024) and 3D-based (Ze et al. 2024) policy models and the proposed FlowPolicy. We evaluate 37 tasks from Adroit and Metaworld across 3 random seeds and report the success rate (%) with standard deviation. $\ast$’ indicates that the NFE of Adaflow is not fixed.
@@
FlowPolicy applies 3D point clouds as a guide to generate execution-specific policies given its rich spatial information. Specifically, we first acquire single-view depth-observation images and then convert the depth information into a point cloud via camera intrinsics and extrinsic. The single-view image is more practically applicable and the point cloud provides a richer spatial representation. Further, a farthest point sampling (FPS) (Qi et al. 2017) is applied to downsample the point cloud to preserve spatial wholeness and reduce redundancy. Finally, we employ a lightweight MLP to encode the point cloud into a compact 3D representation as the visual conditional representation for flow-based policy. With the conditional constraints of visual features, FlowPolicy enables the efficient capture of spatial information that is critical to fine robotic manipulation tasks.
@@
To train a high-quality one-step generator, we employ a two-segment Consistency-FM and use backpropagation to minimize the following equation:
@@
By training the two-segment loss above, two consistent vector fields can be learned, which in turn constructs a two-segment linear trajectory that transforms the noisy trajectory into the target trajectory. In the inference process, the noise trajectories are sampled from the source distribution $a_{0}\in a^{\text{src}}$ and then predicted by a flow model to obtain action trajectories $a_{1}\in a^{\text{tar}}$.
FlowPolicy provides a framework for generating high-quality actions in a one-step process that promotes a trade-off between efficiency and effectiveness.
@@
We choose two preeminent environmental simulators, Adroit (Rajeswaran et al. 2018) and Metaworld (Yu et al. 2021), as our benchmarks for a comprehensive set of 37 tasks. The Adroit benchmark presents three challenges of dexterous manipulation tasks, where the execution of complex tasks is facilitated by the intricate control of a high-dimensional, articulated hand. In contrast, the Metaworld benchmark offers a diverse array of tasks that span the spectrum of difficulty levels, from easy to very hard, typically surmounted with the aid of a two-fingered robotic arm. For the procurement of expert demonstrations, we have harnessed a well-trained heuristic strategy, designed to yield datasets of exceptional quality.
@@
The main objective of this work is to leverage 3D point clouds as visual conditional representations, aiming to elevate the robot manipulation to the real-time level via a consistency flow matching strategy. Therefore, we focus on comparing DP3, a state-of-the-art 3D conditional diffusion-based strategy model. In addition, its lightweight version, simple DP3, is also compared. We also compared state-of-the-art 2D-based approaches, including diffusion policy (DP) (Chi et al. 2023), Adaflow (Hu et al. 2024), and Consistency Policy (CP) (Prasad et al. 2024) and collected the same viewpoint images for them.
@@
We evaluate 20 episodes per 200 training epochs and record the mean of the five highest success rates. Each task is run repeatedly under three different random seeds, and their means and variances as well as inference times are calculated. In addition, the Number of Function Evaluations (NFE) is evaluated as a metric of inference efficiency. The higher the success rate, the better, and the lower the NFE and inference time, the better.
@@
For each task, we generate a compact conditional embedding consisting of a 3D point cloud and the robot’s state with a dimension of 64. Additionally, our model employs an observation horizon of two steps, signifying that it leverages the point clouds from the two most recent time frames as conditional inputs for policy generation. We assign ten expert demonstrations to each task. The image size is cropped to a resolution of 84$\times$84 pixels, and the point cloud is downsampled to 512 points for Adroit or 1024 points for Metaworld using FPS. During the training phase, the weights are updated using the AdamW optimizer with a learning rate of 1e-4 and a batch size of 128. The training process is iterated for 3,000 epochs. EMA decay rate is set to the default value of 0.95. Additionally, both action and state inputs are normalized to the interval [-1, 1] to facilitate the training process, with actions being unnormalized before execution. Our model and all compared methods are developed based on the PyTorch framework and evaluated on a single NVIDIA RTX 2080Ti GPU.
@@
Table 1 reports the average inference time of each model for 3 dexterity tasks from Adroit and 34 different difficulty tasks from Metaworld. Our FlowPolicy demonstrates remarkable results by reducing the average inference time to 19.9ms, improving by more than 7$\times$ (19.9ms vs. 145.7ms) compared to the DP3 model. Even for the lightweight version of Simple DP3, FlowPolicy enhances the inference efficiency by over 3$\times$ (19.9ms vs. 63.0ms). The consistent velocity field provides a more direct linear path, enabling faster and more efficient action generation. Notably, our approach exhibits less variation in inference time. Specifically, FlowPolicy has less fluctuation in the average runtime for both individual and overall tasks, with a potential explanation attributed to the deterministic sampling policy of consistent flow matching. The results show that faster ODE trajectories can be determined and the inference efficiency can be improved by directly defining a straight-line flow, which is crucial for realizing real-time robot operation.